### v1

@@ 4.2 Beyond the Finite Case
In this section we extend the proof of Theorem 1 for finite $T$ to the countable case by a limiting argument, and also prove Proposition 1. We first prove an auxiliary result about absolute integrability of $|X_{Z}|$, which will be used later to apply dominated convergence.
@@
It is interesting to compare Lemma 6 with the classical observation that for symmetric processes, $\mathbb{E}[\sup_{t\in T}|X_{t}|]<\infty$ is equivalent to $\mathbb{E}[\sup_{t\in T}X_{t}]<\infty$, both equivalent to boundedness of the process in the Gaussian case; see (Talagrand,, 1987, p2)(Talagrand,, 2014, Exercise 2.2.2.) and the references therein.
@@
We first prove (78). Clearly, the natural metric for the new process satisfies $\bar{d}((t,s),(t^{\prime},s))=d(t,t^{\prime})$, for any $s\in\{-1,+1\}$ and $t,t^{\prime}\in T$. Now suppose that $P_{\bar{Z}}=P_{Z,S}$ is an arbitrary distribution on $\bar{T}$. We can assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. By the triangle inequality we can show that $\bar{d}(\bar{Z},(t_{0},+1))\leq d(Z,t_{0})+\|2X_{t_{0}}\|_{\Psi_{2}}=d(Z,t_{0})$, implying that $\bar{\sigma}_{\rm m}^{2}\leq\mathbb{E}[\bar{d}^{2}(\bar{Z},(t_{0},+1))]<\sigma_{\rm m}^{2}$. In addition, suppose that $U$ is any random variable such that $R_{P_{Z}}(\sigma^{2})=I(U;Z)$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$ for some $\sigma^{2}$. If we set $\bar{U}=(U,S)$, we have $\mathbb{E}[\bar{d}^{2}(\bar{Z},\bar{U})]=\mathbb{E}[d^{2}(Z,U)]\leq\sigma^{2}$, and moreover, the rate-distortion function on $\bar{T}$ satisfies $\bar{R}_{P_{\bar{Z}}}(\sigma^{2})\leq I(\bar{U};\bar{Z})=I(U,S;Z,S)=I(U;Z)+I(S;Z|U)+H(S|Z)\leq R(\sigma^{2})+2\ln 2$. Then (78) follows by the sub-additivity of the square root function.
@@
$\leq 48\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}_{k}}}(\sigma^{2})}d\sigma$ (88)
$\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}d\sigma+\sqrt{2\ln 2}\sigma_{\rm m},$ (89)
@@
Proof of Theorem 1.
At this point, it only remains to extend the result of Theorem 1 from the finite case to the countable case. We can assume that the integral on the right side of (33) is finite, since otherwise there is nothing to prove. Again, assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}:=\sigma_{\rm m}(P_{Z})=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. Construct $Z_{k}$ as in the proof of Lemma 6. Lemma 6 implies $\mathbb{E}[|X_{Z}|]<\infty$, and since $|X_{Z_{k}}|\leq|X_{Z}|\vee|X_{t_{0}}|$, dominated convergence implies
@@
Then the claim of the lemma follows by (90), (91) and applying the result of Theorem 1 in the finite setting (proved in Section 4.1).
Proof of Proposition 1.
Let $T_{0}$ be the countable subset in the definition of separability. For any $\epsilon>0$, let $Z_{\epsilon}\in T_{0}$ be a random variable such that $|X_{Z}-X_{Z_{\epsilon}}|\leq\epsilon$ and $d(Z,Z_{\epsilon})\leq\epsilon$, which is possible by the separability and continuity assumption. Then $\mathbb{E}[X_{Z}]\leq\mathbb{E}[X_{Z_{\epsilon}}]+\epsilon$, and $W_{\infty}(\mu,\mu^{\prime})\leq\epsilon$, if $\mu^{\prime}$ is the distribution of $Z_{\epsilon}$. Therefore the claim follows by the result for the countable case and taking $\epsilon\downarrow 0$. ∎
5 Sharpness of the Rate-Distortion Integral
We provide several proofs of Theorem 3 (up to constant factors) in this section. The proof in Section 5.3 gives the constant stated in Theorem 3.
5.1 Proof via the Majorizing Measure Theorem
We first give a proof of the lower bound direction assuming a version of the majorizing measure theorem, i.e. the $\delta_{2}$ functional defined in (15) lower bounds the supremum of a Gaussian process. It is relatively easy (compared to the proof of the majorizing measure theorem) to show the equivalence of $\delta_{2}$ and $\gamma_{2}$ up to universal constants, and the direction of $\gamma_{2}\gtrsim\delta_{2}$ can be found in (Talagrand,, 2014, Proposition 3.3.1 & Proposition 3.4.1).
We first consider the case where $n:=|T|<\infty$. Recall the representation of $X_{z}$ as an inner product in (55), which embeds $Y,Z$ in $\mathbb{R}^{n}$, where $Y$ is a normal vector, and recall the typical set $\mathcal{C}$ defined in (54). Define
$\mathcal{D}:=\{y^{N}\in\mathbb{R}^{nN}\colon W_{2}(P_{Y},\widehat{P}_{y^{N}})\leq N^{-\frac{1}{2n}}\}$ (92)
Then by the Wasserstein law of large numbers, $\mathbb{P}[Y^{N}\in\mathcal{D}]=1-o(1)$.
Given $P_{Z}=\mu$, let $P_{YZ}$ be the optimal coupling of $P_{Y}$ and $\mu$ maximizing $\mathbb{E}[\left<Y,Z\right$>] (equivalently, the optimal transport map). Then
$\mathbb{E}[X_{Z}]$ $\geq\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]$ (93)
$=\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>] (94)
$\gtrsim\lim_{N\to\infty}\frac{1}{N}\delta_{2}(\mathcal{C})$ (95)
$\geq\lim_{N\to\infty}\frac{1}{N}\inf_{t\in\mathcal{C}}I_{\mu^{\otimes N}}(t).$ (96)
The above steps are justified below: to show (93), note that for any $y^{N}\in\mathcal{D}$, we have
$\lim_{N\to\infty}\frac{1}{N}\sup_{z^{N}\in\mathcal{C}}\left<y^{N},z^{N}\right$> $\leq\lim_{\epsilon\to 0}\sup_{P_{Y^{\prime}Z^{\prime}}}\mathbb{E}[\left<Y^{\prime},Z^{\prime}\right$>] (97)
$=\lim_{\epsilon\to 0}\sup_{P_{Y^{\prime}Z^{\prime}}}\frac{\mathbb{E}\|Y^{\prime}\|_{2}^{2}+\mathbb{E}\|Z^{\prime}\|_{2}^{2}-W_{2}^{2}(P_{Y^{\prime}},P_{Z^{\prime}})}{2}$ (98)
$=\frac{\mathbb{E}\|Y\|_{2}^{2}+\mathbb{E}\|Z\|_{2}^{2}-W_{2}^{2}(P_{Y},P_{Z})}{2}$ (99)
$=\mathbb{E}[\left<Y,Z\right$>] (100)
where in (97) the sup is over $P_{Y^{\prime}Z^{\prime}}$ for which $W_{2}(P_{Y},P_{Y^{\prime}})\leq\epsilon$ and $W_{2}(P_{Z},P_{Z^{\prime}})\leq\epsilon$. This shows that the right side of (93) is upper bounded by $\mathbb{E}[\left<Y,Z\right>]\cdot\lim_{N\to\infty}\mathbb{E}[1\{Y^{N}\in\mathcal{D}\}]=\mathbb{E}[\left<Y,Z\right$>]. The step (94) follows from $\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}^{c}}]=o(N)$, whose proof is the same as the proof that the second expectation in (58) is $o(N)$. (95) follows since $\delta_{2}(\mathcal{C})\asymp\gamma_{2}(\mathcal{C})$ lower bounds (94) according to the majorizing measure theorem (Talagrand,, 2014). We see (96) by specifying the supremum in (15) as $\mu^{\otimes N}$.
Using the estimate in Lemma 4, we obtain that for any $t\in\mathcal{C}$,
$\liminf_{N\to\infty}\frac{1}{N}I_{\mu^{\otimes N}}(t)\gtrsim\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (101)
by using Fatou’s lemma. This proves Theorem 3 for finite $T$ (except for estimating the constant factor).
@@ 5.1 Proof via the Majorizing Measure Theorem
Moreover, if $P_{U_{k}|Z_{k}}$ achieves the infimum in the definition of $R_{\mu_{k}}(\sigma^{2})$, we can construct $P_{U|Z}$ by setting $P_{U|Z}(|z)=P_{U_{k}|Z_{k}}(|z)$ for $z\in T_{k}$ and $U=t_{0}$ if $z\notin T_{k}$. Assume without loss of generality that $\mathbb{E}_{\mu}[d^{2}(t_{0},Z)]=\sigma_{\rm m}^{2}$. Then $\mathbb{E}[d^{2}(U,Z)]=\sigma^{2}\mu(T_{k})+\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]$. Since $\lim_{k\to\infty}\mathbb{P}[T_{k}^{c}]=0$ and $\mathbb{E}[d^{2}(t_{0},Z)]<\infty$, a basic result from measure theory (which can be proved by dominated convergence) states that $\lim_{k\to\infty}\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]=0$. In particular, for any given $\sigma>0$, there exists $k(\sigma)>0$ such that for all $k>k(\sigma)$, this construction of $P_{U|Z}$ ensures that $\mathbb{E}[d^{2}(U,Z)]<2\sigma^{2}$. Moreover, if $E_{k}:=1\{Z\in T_{k}\}$, we have $I(U;Z)\leq I(U;Z|E)+H(E)=R_{\mu_{k}}(\sigma^{2})+h(\mu(T_{k}))$, where $h(\cdot)$ is the binary entropy function. Taking $k\to\infty$ we have $\mu(T_{k})\to 1$ and obtain
$\lim_{k\to\infty}R_{\mu_{k}}(\sigma^{2})\geq R_{\mu}(2\sigma^{2})$ (103)

### v3

@@ Appendix B Extension of the rate-distortion integral beyond finite $T$
We first prove an auxiliary result about absolute integrability of $|X_{Z}|$, which will be used later to apply dominated convergence.
@@
Lemma 13 may be compared with the classical observation that for symmetric processes, $\mathbb{E}[\sup_{t\in T}|X_{t}|]<\infty$ is equivalent to $\mathbb{E}[\sup_{t\in T}X_{t}]<\infty$, both equivalent to boundedness of the process in the Gaussian case; see [37, p2], [39, Exercise 2.2.2.] and the references therein.
@@
We first prove (131). The natural metric for the new process satisfies $\bar{d}((t,s),(t^{\prime},s))=d(t,t^{\prime})$, for any $s\in\{-1,+1\}$ and $t,t^{\prime}\in T$. Now suppose that $P_{\bar{Z}}=P_{Z,S}$ is an arbitrary distribution on $\bar{T}$. We can assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. By the triangle inequality we can show that $\bar{d}(\bar{Z},(t_{0},+1))\leq d(Z,t_{0})+\|2X_{t_{0}}\|_{\Psi_{2}}=d(Z,t_{0})$, implying that $\bar{\sigma}_{\rm m}^{2}\leq\mathbb{E}[\bar{d}^{2}(\bar{Z},(t_{0},+1))]<\sigma_{\rm m}^{2}$. In addition, suppose that $U$ is any random variable such that $R_{P_{Z}}(\sigma^{2})=I(U;Z)$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$ for some $\sigma^{2}$. If we set $\bar{U}=(U,S)$, we have $\mathbb{E}[\bar{d}^{2}(\bar{Z},\bar{U})]=\mathbb{E}[d^{2}(Z,U)]\leq\sigma^{2}$, and moreover, the rate-distortion function on $\bar{T}$ satisfies $\bar{R}_{P_{\bar{Z}}}(\sigma^{2})\leq I(\bar{U};\bar{Z})=I(U,S;Z,S)=I(U;Z)+I(S;Z|U)+H(S|Z)\leq R(\sigma^{2})+2\ln 2$. Then (131) follows by the sub-additivity of the square root function.
@@
$\lesssim\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}_{k}}}(\sigma^{2})}\,d\sigma$ (141)
$\lesssim\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}\,d\sigma+\sqrt{2\ln 2}\sigma_{\rm m},$ (142)
@@
B.2 Extension of Theorem 1 beyond finite $T$
We extend the result of Theorem 1 from the finite case to the countable case. We can assume that $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}\,d\sigma<\infty$, since otherwise there is nothing to prove. Again, assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}:=\sigma_{\rm m}(P_{Z})=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. Construct $Z_{k}$ and define $E$ as in the proof of Lemma 13. Lemma 13 implies $\mathbb{E}[|X_{Z}|]<\infty$, and since $|X_{Z_{k}}|\leq|X_{Z}|\vee|X_{t_{0}}|$, dominated convergence implies
@@ B.2 Extension of Theorem 1 beyond finite $T$
Then the claim of the lemma follows by (143), (144) and applying the result of Theorem 1 in the finite case (already proved in the main text).
B.3 Proof of Proposition 1
Let $T_{0}$ be the countable subset in the definition of separability. For any $\epsilon>0$, let $Z_{\epsilon}\in T_{0}$ be a random variable such that $|X_{Z}-X_{Z_{\epsilon}}|\leq\epsilon$ and $d(Z,Z_{\epsilon})\leq\epsilon$, which is possible by the separability and continuity assumption. Then $\mathbb{E}[X_{Z}]\leq\mathbb{E}[X_{Z_{\epsilon}}]+\epsilon$, and $W_{\infty}(\mu,\mu^{\prime})\leq\epsilon$, if $\mu^{\prime}$ is the distribution of $Z_{\epsilon}$. Therefore the claim follows by the result for the countable case and taking $\epsilon\downarrow 0$.
B.4 Extension of Theorem 3 beyond finite $T$
@@
Moreover, if $P_{U_{k}|Z_{k}}$ achieves the infimum in the definition of $R_{\mu_{k}}(\sigma^{2})$, we can construct $P_{U|Z}$ by setting $P_{U|Z}(|z)=P_{U_{k}|Z_{k}}(|z)$ for $z\in T_{k}$ and $U=t_{0}$ if $z\notin T_{k}$. Assume without loss of generality that $\mathbb{E}_{\mu}[d^{2}(t_{0},Z)]=\sigma_{\rm m}^{2}$. Then $\mathbb{E}[d^{2}(U,Z)]=\sigma^{2}\mu(T_{k})+\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]$. Since $\lim_{k\to\infty}\mathbb{P}[T_{k}^{c}]=0$ and $\mathbb{E}[d^{2}(t_{0},Z)]<\infty$, a basic result from measure theory (which can be proved by dominated convergence) states that $\lim_{k\to\infty}\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]=0$. In particular, for any given $\sigma>0$, there exists $k(\sigma)>0$ such that for all $k>k(\sigma)$, this construction of $P_{U|Z}$ ensures that $\mathbb{E}[d^{2}(U,Z)]<(1+a)\sigma^{2}$, where $a>0$ is an arbitrary constant. Moreover, if $E_{k}:=1\{Z\in T_{k}\}$, we have $I(U;Z)\leq I(U;Z|E)+H(E)=R_{\mu_{k}}(\sigma^{2})+h(\mu(T_{k}))$, where $h(\cdot)$ is the binary entropy function. Taking $k\to\infty$ we have $\mu(T_{k})\to 1$ and obtain
$\lim_{k\to\infty}R_{\mu_{k}}(\sigma^{2})\geq R_{\mu}((1+a)\sigma^{2})$ (146)